OpenStudy (anonymous):

Find an equation of the line. Write the equation in the standard form. Through (4,3); parallel to 7x-y=7 (Simplify your answer. Type your answer in standard form)

OpenStudy (anonymous):

@johnweldon1993 :)

OpenStudy (johnweldon1993):

Your turn for this one :P So we have the line \[\large 7x - y = 7\] How can you make that look like \(\large y = mx + b\) ?

OpenStudy (anonymous):

7=7x-y

OpenStudy (johnweldon1993):

We want the 'y' by itself on one side of that '=' sign...so lets try adding y to both sides of the equation...and then subtracting 7 from both sides \[\large 7x - y = 7\] afterwe add y to both sides we have \[\large 7x = 7 + y\] and after we subtract 7 from both sides we have \[\large y = 7x - 7\] make sense there??

OpenStudy (anonymous):

yep

OpenStudy (johnweldon1993):

So again...parallel lines...what will be the slope of the line?

OpenStudy (anonymous):

the same slope

OpenStudy (johnweldon1993):

Which is?

OpenStudy (johnweldon1993):

Remember it is 'm' in the equation y = mx + b so it looks like the slope is 7 correct?

OpenStudy (anonymous):

parrallel lines have the same slope

OpenStudy (anonymous):

correct

OpenStudy (johnweldon1993):

So now we know the slope...and we have the point again...so \[\large y - y_0 = m(x - x_0)\] plug in the \(\large x_0 = 4\) and \(\large y_0 = 3\) and also m = 7

OpenStudy (johnweldon1993):

\[\large y - 3 = 7(x - 4)\] so distribute the 7 \[\large y - 3 = 7x - 28\] and add 3 to both sides of the equation \[\large y = 7x - 25\] did all that make sense? :)

OpenStudy (anonymous):

okay

OpenStudy (johnweldon1993):

That didnt sound too reassuring lol whats confusing you there? :)

OpenStudy (anonymous):

No I actually do get it for the first time now haha you're really good at explaining things!

OpenStudy (johnweldon1993):

Oh good ^_^ haha I'm glad then :) anything else you need help with hun?

OpenStudy (anonymous):

nope thats all for now thanks again you're awesome dude

OpenStudy (johnweldon1993):

Glad to help :)

OpenStudy (anonymous):

wait a sec its saying that y=7x-25 is wrong?

OpenStudy (johnweldon1993):

Ohhhh oh oh they want it in standard form...my apologies They want it to look like \[\large -7x + y = -25\]

OpenStudy (anonymous):

oh goody it worked kudos :D

OpenStudy (johnweldon1993):

Lol sorry about that :P

OpenStudy (anonymous):

dont even worry bout it